# PADetBench: Towards Benchmarking Physical Attacks against Object Detection

### v1

@@ 3.4 Evaluation and analysis
To rigorously assess the efficacy of physical attacks on object detection systems, we furnish baseline datasets: clean datasets (without perturbations) and those infused with randomized noise (incorporating arbitrary disturbances in -bounded space). This dual-baseline approach sets the stage for a thorough examination.
Quantifying performance entails employing evaluation metrics that consider both detection accuracy and the performance of the attacks. These metrics comprise by several widely adopted indicators, including average recall (AR) and Attack Success Rate (ASR). AR is calculated as the average of the recall rates:
where TP and FN denote the true positive and false negative counts of the detector, respectively. AR signifies the proportion of correctly identified instances relative to the actual total in the dataset, offering an intuitive gauge of how physical attacks degrade the detection capability of a given target. On the other hand, ASR quantifies the effectiveness of the adversarial perturbations, calculated as:
where and denote the average recall rates of the attack and clean datasets, respectively. ASR provides a direct measure of the extent to which the attacks undermine the detector’s confidence in its predictions.
Furthermore, we enhance our codebase by incorporating several ready-to-use explainability visualization tools, facilitating deeper insights into model behavior. These include Grad-CAM [71] for visualizing the regions of input data that contribute most to the model’s prediction, Shapley value [57] to quantify the individual feature contributions, and t-SNE [83] for reducing dimensionality and visualizing high-dimensional data in a more interpretable manner. These additions empower users to conduct comprehensive analyses beyond mere performance evaluation.
@@ 4.1 Experimental setup
Datasets. 1) Overall experiments. We generate overall datasets with 3 objects, 10 weather conditions, 2 altitude angles, 8 azimuth angles, 5 radius values, 3 spawn points, and 20 physical perturbations, i.e., 7200 samples () for each attack method, in which the physical dynamics are strictly aligned and controlled for impartial comparison. 2) Ablation Studies. We conduct in-depth examinations to explore the individual impact of core physical dynamics: weather conditions, venue, camera distance, azimuth angle, altitude angle within a hemispherical space. Accomplishing this involves generating focused sub-benchmarks, each consisting of 100 samples.
Physical attacks. we generate 24 datasets, including 20 physically noised datasets that correspond to 20 physical attacks, an extra 2 clean datasets and 2 randomly noised datasets for comparison of vehicle detection and person detection, respectively.
Object detectors. We evaluate 48 object detectors, as shown in Table 1, by integrating MMDetection [9] into our evaluation pipeline, in which the selected detectors are trained on the COCO [50] dataset.
Therefore, we conduct a total of 8064 () groups of the experiment, which are conducted with NVIDIA Geforce 4090.
@@ 4.2 Overall experiments and analysis
We present the comprehensive results of vehicle detection against physical attacks in Fig. 2. Additionally, Fig. 3 and Fig. 4 show visualized analyses of the experimental results from detection and attack perspectives, respectively. Experimental results of person detection and corresponding detailed numerical results are listed in the Appendix. From these analyses, several key observations emerge:
Detection perspective: 1) Vehicle detection performance is significantly impacted by physical attacks, with the average recall rates of detectors decreasing up to 50%, as shown in Fig. 4. However, pedestrian detection performance is less affected regarding various attacks, with the average recall rates of detectors decreasing by less than 20%, as shown in Fig. 3. The potential reason is that the stronger physical perturbations are optimized with consideration of 3D space and accommodate more complex physical dynamics, while physical attacks aiming to fool person detectors are commonly performed with optimized 2D patches, which work well in particular physical dynamics, as shown in the ablation experiments, which empirically demonstrate the pressing need and necessity of a comprehensive and rigorous benchmark for physical attacks. 2) For different detectors, the performance varies significantly, with some detectors exhibiting superior robustness against physical attacks like EfficientNet, YOLO series and RTMDet of one-stage detectors. Besides, DDQ also shows significant adversarial robustness among transformer-based detectors. While the rest of the detectors show varying lower levels of robustness, the SOTA detection performance is not necessarily indicative of adversarial robustness. Consequently, the benchmark also serves as a robustness indicator.
Attack perspective: 1) For vehicle detection, different physical attacks exhibit varying levels of effectiveness, with some attacks achieving ASR values exceeding 70% like ACTIVE, and others failing to surpass 20%. Most of the physical attacks hard to fool the latest SOTA detectors, such as EfficientNet, YOLO series, and RTMDet. This phenomenon is caused by the victim models of the attack method lagging behind the development of the detection method, which also motivates us to fill this gap. 2) For person detection, the ASR values of physical attacks are generally lower than those for vehicle detection, with the majority of attacks achieving ASR values below 20%, even lower than 0% (see more explanation in the Appendix). The relatively strongest attack method is AdvTexture, which elaborates on a 2D patch but with tricks for 3D space. This also demonstrates the gap between 2D perturbations and 3D physical space.
@@ 4.3 Ablation experiments and analysis
Except for the overall experiments, we also conduct ablation experiments to investigate the impact of physical world factors. We show the results of 3 physical dynamics, including weather, distance, and camera viewing angle, in Fig. 5 and Fig. 6, respectively. More experiments on other dynamics are provided in the Appendix. From these analyses, several key observations emerge: 1) Physical attack performance can be easily swayed by physical dynamics. This phenomenon is consistent with our intuition and emphasizes the importance of strictly aligning physical dynamics when evaluating physical attacks, which are often underestimated by previous works. 2) We also observe a gap between the attack performance of our benchmark and the reported performance in the original papers (supplemented in Appendix). Two reasons may contribute to this gap: the first is the adopted SOTA detectors in our benchmark, which are more robust than the victim models in the original papers, and the second is that our benchmark provides more comprehensive and strict evaluation datasets and physical dynamics, which are more challenging for the attack methods. Please refer to the Appendix for more experiments, detailed analysis and discussion.
@@ 5 Discussion
Limitations. For now, PADetBench mainly focuses on the physical attacks on object detection models. In the future, we will extend the benchmark to evaluate other vision tasks, such as instance segmentation, 3D object detection, and depth estimation.
Potential impacts. 1) Positive impacts: In-depth understanding of the robustness of object detection models against physical attacks, which can guide the development of more robust models. 2) Negative impacts: The benchmark may be misused to conduct physical attacks in real life, which may pose a threat to the security-critical applications involving intelligent visual perception.
@@ 6 Conclusion
In conclusion, this work aims to meet the critical need for a standardized and rigorous benchmark for physical attacks on object detection models. Our benchmark leverages the CARLA simulator for consistent physical dynamics and includes 20 attack methods, 48 detectors, comprehensive dynamics, and metrics from different perspectives. It also provides end-to-end pipelines for dataset generation, attack and detection evaluation, and experimental results analysis. Our extensive evaluation and analyses provide valuable insights for future research directions. The supporting codebase is publicly available and we will continuously supplement it, aiding the development of physically robust object detection methods.
7 Acknowledgment
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust.
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes]
Did you discuss any potential negative societal impacts of your work? [Yes]
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Please refer to the GitHub repository: https://github.com/JiaweiLian/Benchmarking_Physical_Attack for the code and instructions.
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A] This benchmark does not involve a training process.
Did you report error bars (e.g., to the random seed after running experiments multiple times)? [N/A]
Did you include the total amount of computing and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
- (a)
Did you mention the license of the assets? [Yes] Please refer to the Appendix for the license.
Did you include any new assets either in the supplemental material or as a URL? [Yes] Please refer to the GitHub repository (https://github.com/JiaweiLian/Benchmarking_Physical_Attack) for details.
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ A.1 Mini-test
Town4 | A small town embedded in the mountains with a special "figure of 8" infinite highway. |
Town5 | Squared-grid town with cross junctions and a bridge. |
It has multiple lanes per direction. Useful to perform lane changes. | |
@@
Town8 | Secret "unseen" town used for the Leaderboard challenge. |
Town9 | Secret "unseen" town used for the Leaderboard challenge. |
@@
Town11 | A Large Map that is undecorated. |
Serves as a proof of concept for the Large Maps feature. | |
Town12 | A Large Map with numerous different regions, |
including high-rise, residential and rural environments. |
Full list of the optional maps, where Town8 and Town9 are unseen for competition. Please refer to CARLA [17] documentary for more details.
A.2 Corresponding config files of the selected detectors
The corresponding config files of the selected detectors are listed in Table 3.

### v3

@@ 3.4 Evaluation and Analysis
Evaluation metrics. To rigorously assess the efficacy of physical attacks on object detection systems, we furnish baseline datasets: clean datasets (without perturbations) and those infused with randomized noise (incorporating arbitrary disturbances in -bounded space). This dual-baseline approach sets the stage for a thorough and fair examination. Quantifying performance entails employing evaluation metrics that consider object detection and adversarial attack performance. These metrics comprise several widely adopted indicators, including mean average precision (mAP), mean average recall (mAR), and attack success rate (ASR). mAP and mAR are calculated as the mean value of average precisions and recalls at recall and precision levels over classes, respectively, i.e., and . Precision rate and recall rate are calculated as and , respectively, where TP, FP, and FN denote the true positive, false positive, and false negative counts of the detector, respectively. On the other hand, ASR quantifies the effectiveness of the adversarial perturbations, calculated as , where and denote the value of adopted metric on the attack and clean datasets, respectively. ASR directly measures the extent to which the attacks undermine the detector’s performance.
Advocacy of mAR for physical attacks. Adversarial attacks aim to induce mispredictions, i.e., to maximize error rate, which is the mathematical expectation of incorrect predictions written as:
where is 1 for a correct prediction and 0 otherwise, and and represent the ground truths and predicted results of all objects, respectively. According to the calculation of performance metrics for detection, we can rewrite the error rate as follows:
Therefore, mAR is a more direct and intuitive metric for evaluating the effectiveness of physical attacks on object detection models. We use mAR as the primary metric in the main manuscript, while mAP is also provided for reference.
Evaluation perspectives. Specifically, we use mAP50, i.e., the confidence threshold of 0.5, to evaluate the overall performance of object detection, which is widely adopted in the object detection community. mAR50 is adopted to signify the proportion of correctly identified instances relative to the actual total in the dataset, offering an intuitive gauge of how physical attacks degrade the detection capability of a given adversarial target. However, mAR50 and mAP50 cannot fully reflect the performance of object detection models, especially when the confidence score of an adversarial object is significantly dropped but still higher than the threshold. To address this issue, we also use mAR50:95 and mAP50:95, calculated as the mean value over the range of 0.5 to 0.95 of the confidence threshold, to provide a more comprehensive evaluation of the object detection models. From the perspective of physical attacks, we use ASR over the detection metrics mAP50, mAR50, mAP50:95, and mAR50:95 to evaluate the effectiveness of physical attacks on object detection models, ensuring a comprehensive and impartial assessment. Moreover, we also visualize the distribution of evaluation performance using violin plots, which can provide a more intuitive understanding of the performance of object detection models and physical attacks, respectively.
Analysis tools. Furthermore, we enhance our codebase by incorporating several ready-to-use explainability visualization tools, facilitating more profound insights into model behavior. These include Grad-CAM [132] for visualizing the regions of input data that contribute most to the model’s prediction, Shapley value [133] to quantify the individual feature contributions, and t-SNE [134] for reducing dimensionality and visualizing high-dimensional data in a more interpretable manner. These additions empower users to conduct comprehensive analyses beyond mere performance evaluation.
@@ 4.1 Experimental setup
Datasets. 1) Overall experiments. We generate overall datasets with three objects, ten weather conditions, two altitude angles, eight azimuth angles, five radius values, three spawn points, and 23 physical perturbations, i.e., 7200 samples () for each attack method, in which the physical dynamics are strictly aligned and controlled for impartial comparison as shown in Fig. 3. Please note that these parameters are adjustable in the pipeline, and the datasets can be easily generated with different settings as needed. 2) Ablation Studies. We conduct in-depth examinations to explore the individual impact of core physical dynamics: weather conditions, venue, camera distance, azimuth angle, and altitude angle within a hemispherical space. Accomplishing this involves generating focused sub-benchmarks, each consisting of 100 samples.
Physical attacks. We generate 24 datasets for comprehensive evaluation, including 20 physically noised datasets corresponding to 20 physical attacks, an extra two clean datasets, and two randomly noised datasets for comparison of vehicle and person detection, respectively. To evaluate the attack transferability, we also adopt perturbations optimized for aerial detection [135] and depth estimation [6, 87] in the experiments. Furthermore, we generate four extra datasets concerning traffic sign detection to show the easy extension of the benchmark to other objects (refer to A.2 for more details). The involved physical attacks are detailed in Table 1.
Object detectors. We evaluate 48 object detectors covering mainstream types, such as one- and two-stage and transformer-based detectors, as shown in Table 2, by integrating MMDetection [131] into our evaluation pipeline.
Therefore, we conduct 8256 () groups of the experiment, which are conducted with NVIDIA GeForce 4090.
@@ 4.2 Overall experiments and analysis
We present the comprehensive results of vehicle detection against physical attacks in Table 3 and Figure 6. Additionally, Figures 7 and 8 provide visualized analyses of the experimental results from a detection perspective, while Figures 9 and 10 present the results from an attack perspective. More experimental results and corresponding detailed numerical data are listed in B. From these evaluations, several vital observations emerge.
1) Detection Perspective.
Vehicle detection performance is significantly impacted by physical attacks, with the average recall rates of detectors decreasing by up to 50%, as shown in Figure 8. In contrast, pedestrian detection performance is less affected by various attacks, with the average recall rates of detectors decreasing by less than 20%, as shown in Figure 7. The potential reason for this difference is that stronger physical perturbations targeting vehicles are optimized considering 3D space and accommodate more complex physical dynamics. On the other hand, physical attacks aimed at fooling person detectors are commonly performed using optimized 2D patches, which are effective under specific physical dynamics. This observation is further detailed in the ablation experiments (B.2.2), which empirically demonstrate the pressing need for a comprehensive and rigorous benchmark for physical attacks.
The performance of different detectors varies significantly. Some detectors exhibit superior robustness against physical attacks, such as EfficientNet, the YOLO series, and RTMDet among one-stage detectors. Additionally, DDQ demonstrates notable adversarial robustness among transformer-based detectors. Other detectors show varying degrees of lower robustness, and interestingly, state-of-the-art detection performance does not necessarily correlate with adversarial robustness. Consequently, the benchmark also serves as an essential indicator of robustness.
2) Attack perspective.
For vehicle detection, different physical attacks exhibit varying levels of effectiveness. Some attacks achieve ASR values exceeding 70%, such as ACTIVE, while others fail to surpass 20%. Most physical attacks struggle to fool the latest state-of-the-art (SOTA) detectors, including EfficientNet, the YOLO series, and RTMDet. This phenomenon is attributed to the attack methods often lagging behind the rapid development of detection techniques, which motivates us to bridge this gap.
For person detection, the ASR values of physical attacks are generally lower than those for vehicle detection, with most attacks achieving ASR values below 20%. The most effective attack method is AdvTexture, which uses a 2D patch with additional techniques to account for 3D space. This phenomenon highlights the significant gap between 2D perturbations and 3D physical environments, emphasizing the challenges in effectively transferring adversarial attacks from controlled 2D settings to more complex 3D scenarios. Furthermore, it underscores the need to develop more sophisticated attack strategies that can better account for the intricacies of 3D physical dynamics.
@@ 4.3 Ablation experiments and analysis
In addition to the overall experiments, we also conduct ablation experiments to investigate the impact of physical world factors. We show the results of 3 physical dynamics, including weather, distance, and camera viewing angle, in Fig. 9 and Fig. 10, respectively. More experiments on other dynamics are provided in B.3 and B.5. From these evaluations, several key observations emerge:
We also observe a gap between the ablation attack performance of our benchmark and the reported performance in the original papers (refer to B.2.1 for more details). Two reasons may contribute to this gap: the first is the adopted SOTA detectors in our benchmark, which are more robust than the victim models in the original papers, and the second is that our benchmark provides more comprehensive and strict evaluation datasets and physical dynamics, which are more challenging for the attack methods.
These observations empirically demonstrate the pressing need for a comprehensive and rigorous benchmark for physical attacks. Please refer to B for more experiments, detailed analysis, and discussion.
@@ 5 Discussion
5.1 Where are we?
1) Lack of alignment and comprehensiveness in physical dynamics.
Existing works often suffer from limitations in comprehensiveness and fail to strictly control and align physical dynamics, as illustrated in Fig. 2. Previous studies, such as those by Zhong et al. [85] and Dong et al. [66], have demonstrated that physical dynamics can be exploited to deceive DNNs. This observation underscores the critical necessity of aligning these dynamics to ensure fair and accurate evaluations. Without a comprehensive and rigorously aligned study, researchers cannot accurately gauge the progress in this research domain, which impedes advancements in the field. A robust benchmark that meticulously controls and aligns physical dynamics is essential for fostering the development of more resilient and reliable object detection models in the face of physical adversarial attacks.
2) Discrete and naive physical adaptation.
While theoretically, well-studied digital attacks should enhance the effectiveness of physical attacks, the practical outcomes often fall short of expectations. This discrepancy primarily stems from the fact that theoretical advancements do not always survive the cross-domain transformations ( as discussed in Section 2.2. These transformations are essential for translating digital models into real-world applications, but introduce significant challenges that are not adequately addressed by existing methods. Current research typically employs discrete and simplistic augmentation techniques to simulate physical dynamics. However, these methods fail to capture the real-world physical scenarios’ continuous and complex nature. As a result, the performance gaps observed in our ablation experiments (see B.2.1) are not surprising. These gaps highlight the limitations of current approaches and emphasize the need for a more comprehensive and rigorous benchmark.
5.2 Where to go?
1) Comprehensive and Physically Aligned Benchmark
A comprehensive and physically aligned benchmark is essential for evaluating physical attacks on object detection models. Such a benchmark ensures rigorous and unbiased assessments, enabling researchers to identify the strengths and weaknesses of various attack methods and defense mechanisms. Providing detailed insights facilitates the development of more robust and resilient object detection models. This move is crucial for enhancing the security and reliability of AI systems in real-world applications, where the ability to withstand physical attacks is paramount.
A well-designed benchmark should encompass various physical scenarios, including lighting conditions, angles, and environmental factors. It should also include diverse datasets and evaluation metrics to ensure the models are tested under realistic conditions. By doing so, the benchmark can help identify specific areas where current models and defenses fall short, guiding future research and development efforts. Ultimately, a comprehensive and physically aligned benchmark will drive innovation and improve AI systems’ overall security and performance in practical settings.
2) Rigorous and Differentiable Modeling of Cross-Domain Transformations
Accurate modeling of cross-domain transformations is crucial for both physical attacks and defenses. These transformations involve converting digital perturbations into physical perturbations and vice versa, which introduces significant challenges due to the differences in the domains. While existing works have attempted to use differentiable neural renderers to generate adversarial examples automatically, these methods often have limited modeling capabilities and struggle to align physical factors between perturbed and clean images accurately.
The advent of large foundation models offers a promising direction for addressing these limitations. Large-scale data and advanced foundation models can provide the necessary complexity and detail to more rigorously model physical dynamics. By leveraging these resources, researchers can develop more sophisticated methods for simulating real-world physical perturbations, leading to more effective attack simulations and defensive strategies.
For example, large foundation models can be trained on extensive datasets that capture various physical environments and conditions. This training can help the models learn to account for various physical factors, such as lighting variations, texture properties, and camera distortions. Differentiable rendering techniques can also be integrated with these models to ensure the generated perturbations are realistic and physically plausible.
In summary, exploring how to model physical dynamics more rigorously using large-scale data and foundation models is a promising direction. This approach has the potential to significantly improve the realism and effectiveness of both attack simulations and defensive strategies, thereby advancing the field of physical security in AI systems.
@@ 6 Conclusion
In conclusion, we develop a comprehensive simulation-based benchmark to evaluate physical attacks under controlled conditions rigorously. This benchmark encompasses 23 physical attacks, 48 object detectors, and detailed physical dynamics, all supported by end-to-end pipelines. The benchmark is designed to be flexible and scalable, facilitating the easy integration of new attacks, models, and vision tasks. Through extensive evaluations involving over 8,000 tests, we have identified vital algorithm limitations and provided valuable insights. We hope this benchmark will advance research in physical adversarial attacks, fostering the development of more robust and reliable models.
CRediT authorship contribution statement
Jiawei Lian: Conceptualization, Data curation, Formal analysis, Writing - original draft, Writing - review and editing. Jianhong Pan: Data curation, Formal analysis, Writing - original draft. Lefan Wang: Formal analysis, Writing - original draft. Yi Wang, Shaohui Mei, and Lap-Pui Chau: Supervision, Writing - review and editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision, funded by The Hong Kong Jockey Club Charities Trust. This work was supported partly by the National Natural Science Foundation of China (No. 62171381).
Data availability
The datasets used in this study are publicly available and can be accessed through https://github.com/JiaweiLian/PADetBench.
Appendices
A Additional content of the Benchmark
A.1 Mini-test
A.2 The adaptability of the benchmark
A.3 Corresponding config files of the selected detectors
A.4 Explanation about the selected objects
A.5 The necessity of the benchmark
A.5.1 Utilities of the benchmark
A.5.2 Potential applications of the benchmark
A.6 Limitations and potential impacts
B Additional content of the Experiments
B.1 Generated data for ablation studies
B.2 A detailed illustration of the performance gap
B.2.1 Performance gap between the benchmark and the original papers
B.2.2 Performance gap between attacks against vehicle and person detection
B.3 Supplemented experiments analysis and discussion
B.3.1 Detection perspective
B.3.2 Attack perspective
B.4 Additional oversall experimental experiments
B.5 Additional ablation experiments
B.5.1 Ablation study on physical dynamics
B.5.2 Ablation study on training dataset
B.5.3 Ablation study on 2D and 3D perturbations
C User feedback
@@ A.1 Mini-test
Town4 | A small town embedded in the mountains with a special ”figure of 8” infinite highway. |
Town5 | Squared-grid town with cross junctions and a bridge. It has multiple lanes per direction. Useful to perform lane changes. |
@@
Town8 | Secret ”unseen” town used for the Leaderboard challenge. |
Town9 | Secret ”unseen” town used for the Leaderboard challenge. |
@@
Town11 | A Large Map that is undecorated. Serves as a proof of concept for the Large Maps feature. |
Town12 | A Large Map with numerous different regions, including high-rise, residential and rural environments. |
A.2 The adaptability of the benchmark
We provide a detailed illustration of the scene diversity of the benchmark in Table 4 and Fig. 12, where the optional maps are listed with their descriptions. In addition, we display the extendable vehicles, pedestrians, and traffic signs in Fig. 13, Fig. 14, and Fig. 15, respectively, which can be easily extended to evaluate other objects in the benchmark. The users are also allowed to export any customized scenes and objects to the benchmark as needed, which can be easily integrated into the benchmark.
A.3 Corresponding config files of the selected detectors
The corresponding config files of the selected detectors are listed in Table 5. Specifically, 1-25 and 26-40 are CNN-based One-stage and Two-stage object detectors, respectively. 41-48 are Transformer-based object detectors. The corresponding config files of the detectors are available in our codebase or MMDetection [131] toolbox.
@@ A.3 Corresponding config files of the selected detectors
A.4 Explanation about the selected objects
According to a survey [86] published in TPAMI 2024, most physical attacks against object detection are optimized for specific target categories, such as vehicles, persons, and a few for traffic signs. In line with this, we have chosen vehicles and pedestrians as the representative target categories, to evaluate the robustness of object detectors against physical attacks. .
In order to ensure the validity of our benchmark for different types of objects, we have demonstrated that our benchmark can be easily extended to other target categories, as shown by the experiments conducted on traffic sign in Table 20. The benchmark is designed to evaluate the robustness of object detectors against physical attacks in various aligned scenarios for ensuring fairness. It can be extended to other target categories with minimal modifications.
We have thoroughly reviewed over forty physical attack methods, and we found that most of these methods conducted experiments under unaligned conditions and without fair comparisons. This lack of clarity hinders the accurate assessment of the progress of physical adversarial attacks and the development of physical adversarial robustness. Therefore, we are motivated to establish a comprehensive and rigorous benchmark for physical attacks to address these limitations and provide a solid foundation for future research.
A.5 The utility of the benchmark
In this section, we sumarize our motivation and provide the potential applications of the benchmark.
A.5.1 Utilities of the benchmark
Standardization and Fair Evaluation: The primary utility of PADetBench lies in its ability to standardize the evaluation of physical attacks against object detection models. By ensuring that all evaluations are conducted under the same physical dynamics, PADetBench eliminates inconsistencies found in real-world experiments, making it a fair and rigorous benchmark.
Comprehensive Coverage: PADetBench includes 23 physical attack methods and evaluates 48 state-of-the-art object detectors, providing a comprehensive coverage that enables researchers to compare and contrast various models and attack strategies.
A.5.2 Potential applications of the benchmark
Research and Development: Researchers developing robust object detection models or physical attack strategies need a benchmark to evaluate and compare their approaches.
Security Assessments: Security teams need to assess the robustness of deployed object detection systems in critical infrastructure.
Regulatory Compliance: Regulatory bodies require evidence of robustness and security for autonomous systems.
Product Testing: Companies developing autonomous vehicles or security systems need to test their products under various physical attack scenarios.
Educational Purposes: Educators and students need resources to understand the vulnerabilities of object detection models.
A.6 Limitations and potential impacts
Limitations
For now, PADetBench primarily focuses on evaluating the robustness of object detection models against physical attacks. In the future, we plan to extend the benchmark to include other vision tasks, such as instance segmentation, 3D object detection, and depth estimation. This expansion will provide a more comprehensive evaluation framework that covers a broader range of computer vision applications.
Potential Impacts
1) Positive Impacts: The in-depth understanding gained through PADetBench will contribute significantly to the development of more robust object detection models. By identifying vulnerabilities and limitations, researchers and practitioners can design improved algorithms that are better equipped to handle physical adversarial attacks. This enhanced robustness is crucial for real-world applications where reliability and accuracy are paramount.
2) Negative Impacts: While the benchmark provides valuable insights, there is a risk that it could be misused to conduct physical attacks in real-life scenarios. Such misuse could threaten the security of critical applications involving intelligent visual perception systems. Therefore, it is essential to promote responsible use of the benchmark and to emphasize the importance of ethical considerations in research and development.